40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 10161     Entry time: Wed Jul 9 08:50:30 2014
Author: Akhil 
Type: Update 
Category: General 
Subject: Weekly Update 

 

Last Week's Work:

  • Worked on the setup to mitigate timing issues arising due to the non-synchronization of clocks of the Frequency counter and Raspberry Pi .
  • Took measurements with the mentioned setup and generated PSD plots of the FC.
  • Completed the setup for phase measurements by using an external ADC.

 

Work Plan for this Week:

  • Complete the installation of the Mini Circuits Frequency Counter on the EPICS. This involves installation of EPICS base on Raspberry Pi, creating an IOC server on the R Pi and writing the data from the FC into a specific IOC  channel. 
  • Complete phase measurements and obtain the delay in the FC thus completing the characterization of the FC.
  • Install the FC at a suitable place inside the 40m so that the beat note system can be remotely managed from any of the Control computers thus effectively replacing the spectrum analyzer(This will be done with proper supervision once the recently ordered FC is shipped)

 

Inside the 40m :

  • I will be going inside the lab today around 9 am with Manasa to make a plan about where the FC must be placed and the routing of the RF cables and the cables which run  into computers from the FC.
  • Once the channel is created and tested, we will install the FC inside the lab possibly by the end of this week or by next week.
ELOG V3.1.3-